# K-Edit: Language Model Editing with Contextual Knowledge Awareness

### v1

@@ 1 Introduction
The ability to pinpoint and edit knowledge within the parameters of a Large Language Model (LLM) is a powerful tool for maintaining alignment with an ever-changing world. Existing approaches to direct model editing make shallow edits that fail to integrate with contextual knowledge the model possesses (Zhong et al. 2023). This work proposes and explores a simple solution to better enhance knowledge reasoning over edits by including contextual and associated knowledge in the editing process. We use the contextual consistency of knowledge graphs when edited to improve the consistency of language models under direct model editing.
Direct Model Editing (or Knowledge-based Model Editing) aims to change the implicit knowledge in an LLM’s weights with no other changes to how the model is used (Wang et al. 2023b; Mazzia et al. 2023; Yao et al. 2023). Prior methods have focused on the ability to edit and recall a new fact without affecting neighboring facts or general generation ability (Zhu et al. 2020; Mitchell et al. 2021; Meng et al. 2022a, b). Recently, however, researchers have begun to question the effectiveness of these edits through benchmarks evaluating the ability of models to use these edits for multi-hop reasoning (Zhong et al. 2023). For example, if we edit the British Prime Minister to be Rishi Sunak instead of Boris Johnson, we would expect the model to be able to answer Who is the British Prime Minister’s wife? with the correct answer Akshata Murty. Existing direct model editing techniques largely fail in this regard. We address this issue by utilizing the relational information in knowledge graphs.
@@ 2.3 Batch Editing
In order for our approach, K-Edit, to succeed, we rely on the ability to make edits in a batch. Our approach works with any such algorithm. For our experiments, we use MEMIT, a state-of-the-art batch editing algorithm that can scale to thousands of edits at once (Meng et al. 2022b). MEMIT is the best option to use with K-Edit, and the only option that can scale to large numbers of edits.
@@ 3 K-Edit Algorithm
MQuAKE MultiCounterfact
#examples 3,000 10,000
#relations 36 34
Table 1: Dataset statistics.
1-edit 2-edits 3-edits 4-edits
2-hops 513 487 NA NA
3-hops 356 334 310 NA
4-hops 224 246 262 268
@@ 4.1 Datasets
We use two datasets, MQuAKE (Zhong et al. 2023) and MultiCounterFact(Meng et al. 2022b). Dataset statistics can be found in Table 2.
The primary dataset, MQuAKE, is used for evaluating multi-hop reasoning on edits (Zhong et al. 2023). This dataset contains 3,000 multi-hop questions (2-4 hops) with each answer based on 1-4 counterfactual edits. For example, the question “What is the headquarters location for the broadcaster of True Detective?” is paired with two edits: (True Detective, broadcaster, HBO $\rightarrow$ ABC) and (ABC, headquarters location, New York City $\rightarrow$ Los Angeles). These edits are not real in the world, but are plausible counterfactual edits that could be applied. This example was a two-hop question with two edits, however, the number of edits does not need to be the same as the number of hops. The exact breakdown can be found in Table 2.
We also use a subset of MultiCounterFact dataset (Meng et al. 2022b) derived from CounterFact (Meng et al. 2022a). This dataset contains 30,000 counterfactual edits and is used to evaluate the strength of an edit, whether unaffected neighboring relations remain unchanged, and the generation ability of the model post editing. The MultiCounterFact modification ensures there are no direct conflicts between edits in the dataset (i.e. no duplicate edits for a given subject and relation). We use a subset of 10,000 questions for our experiments.
@@ 4.2 Baselines
FT-W is a finetuning approach with weight decay to avoid altering other memories or changing unrelated model behavior (Zhu et al. 2020).
MEND uses a hypernetwork meta-trained to modify fine-tuning gradients so that they produce higher quality edits (Mitchell et al. 2021).
ROME uses counterfactual tracing to locate what layer of the model contains the information to be edited before updating the MLP component in that layer (Meng et al. 2022a).
MEMIT (presented above) uses a similar objective to ROME but modified to enable editing large sets of facts at once, which none of the other baselines are able to do well (Meng et al. 2022b). We consider this to be the main baseline as it is also the editing mechanism we use in our K-Edit implementation.
@@ 6 Related Works
These approaches include aforementioned works (Zhu et al. 2020; Mitchell et al. 2021; Meng et al. 2022b, a). Other works use a hypernetwork to predict parameter updates for a model Cao, Aziz, and Titov (2021)
@@
Outside of direct model editing approaches (Zhu et al. 2020; Meng et al. 2022a; Mitchell et al. 2021; Meng et al. 2022b; Ma et al. 2023), other approaches have sought to adapt to changing knowledge via storing external edit memories and employing various in-context learning and retrieval mechanisms. IKE and others add all edits in the LLM context (Zheng et al. 2023; Onoe et al. 2023); SERAC identifies if a related fact has been edited and retrieves it (Mitchell et al. 2022); MeLLo and DeepEdit iteratively retrieve edited memories during the generation and decoding process (Zhong et al. 2023; Wang et al. 2024).
We note that these retrieval-based editing approaches are fundamentally different than direct model editing approaches. Direct model editing leaves the model architecture and code entirely the same with the only change being to parameter weights. This means any system developed to use the original model can use the updated model with no other changes. In contrast, retrieval-based editing systems fundamentally change how the model is used and functions. For example, MeLLo requires setting up a retrieval database, iteratively prompting the LLM, making multiple calls to the database, and prompting the LLM to create further sub-questions (Zhong et al. 2023). This fundamentally changes the characteristics of the model, increases latency, and alters model behavior for non-QA tasks. An alternate line of work aims to supplement the parametric knowledge of LLMs by augmenting the model prompt with information retrieved from an indexed corpus (Lewis et al. 2020) or Knowledge Graphs (Markowitz et al. 2024).
@@
Other papers have looked into other aspects of contextual knowledge for edits such as reversibility (Ma et al. 2023), multi-linguality (Wang, Haddow, and Birch 2023; Si, Zhang, and Zhang 2024), or other implied logical conditions given an edit (Wang et al. 2023a; Cohen et al. 2023; Li et al. 2023).
@@ Appendix C MEMIT
MEMIT (Meng et al. 2022b) works as follows. MEMIT determines a small change to the $L$-th layer activation of the last token of $s$ (e.g. [Kingdom] in the example above) such that $t_{r}(s)$ produces the desired output $o^{*}$. It then modifies a subset of the model weights to produce the desired activation for a whole batch of such edits.
@@ Appendix E Creating New Relation Templates with LLMs
Figure 2: The above prompt can be used to create relation templates of the form $t_{r}(s)$. These are cloze statements in which the next tokens would match the entity referred to from the subject entity {subject} using relation {relation}. To use this prompt, “{relation}” would be replaced with relation $r$.

### v2

@@ Abstract
K-Edit: Language Model Editing with Contextual Knowledge Awareness
Elan Markowitz,1, 2 Anil Ramakrishna,1 Ninareh Mehrabi,1 Charith Peris,1 Rahul Gupta,1 Kai-Wei Chang,1, 3 Aram Galstyan1 1Amazon AGI, 2University of Southern California - Information Sciences Institute, 3University of California - Los Angeles esmarkow@usc.edu
@@ 1 Introduction
The ability to pinpoint and edit knowledge within the parameters of a Large Language Model (LLM) is a powerful tool for maintaining alignment with an ever-changing world. Existing approaches to direct model editing make shallow edits that fail to integrate with contextual knowledge the model possesses (Zhong et al., 2023). This work proposes and explores a simple solution to better enhance knowledge reasoning over edits by including contextual and associated knowledge in the editing process. We use the contextual consistency of knowledge graphs when edited to improve the consistency of language models under direct model editing.
Direct Model Editing (or Knowledge-based Model Editing) aims to change the implicit knowledge in an LLM’s weights with no other changes to how the model is used (Wang et al., 2023b; Mazzia et al., 2023; Yao et al., 2023). Prior methods have focused on the ability to edit and recall a new fact without affecting neighboring facts or general generation ability (Zhu et al., 2020; Mitchell et al., 2021; Meng et al., 2022a, b). Recently, however, researchers have begun to question the effectiveness of these edits through benchmarks evaluating the ability of models to use these edits for multi-hop reasoning (Zhong et al., 2023). For example, if we edit the British Prime Minister to be Rishi Sunak instead of Boris Johnson, we would expect the model to be able to answer Who is the British Prime Minister’s wife? with the correct answer Akshata Murty. Existing direct model editing techniques largely fail in this regard. We address this issue by utilizing the relational information in knowledge graphs.
@@ 2.3 Batch Editing
In order for our approach, K-Edit, to succeed, we rely on the ability to make edits in a batch. Our approach works with any such algorithm. For our experiments, we use MEMIT, a state-of-the-art batch editing algorithm that can scale to thousands of edits at once (Meng et al., 2022b). MEMIT is the best option to use with K-Edit, and the only option that can scale to large numbers of edits.
@@ 3.2 Scaling Contextual Edits
MQuAKE MultiCounterfact
#examples 3,000 10,000
#relations 36 34
Table 1: Dataset statistics.
1-edit 2-edits 3-edits 4-edits
2-hops 513 487 NA NA
3-hops 356 334 310 NA
4-hops 224 246 262 268
@@ 4.1 Datasets
We use two datasets, MQuAKE (Zhong et al., 2023) and MultiCounterFact(Meng et al., 2022b). Dataset statistics can be found in Table 2.
The primary dataset, MQuAKE, is used for evaluating multi-hop reasoning on edits (Zhong et al., 2023). This dataset contains 3,000 multi-hop questions (2-4 hops) with each answer based on 1-4 counterfactual edits. For example, the question “What is the headquarters location for the broadcaster of True Detective?” is paired with two edits: (True Detective, broadcaster, HBO $\rightarrow$ ABC) and (ABC, headquarters location, New York City $\rightarrow$ Los Angeles). These edits are not real in the world, but are plausible counterfactual edits that could be applied. This example was a two-hop question with two edits, however, the number of edits does not need to be the same as the number of hops. The exact breakdown can be found in Table 2.
We also use a subset of MultiCounterFact dataset (Meng et al., 2022b) derived from CounterFact (Meng et al., 2022a). This dataset contains 30,000 counterfactual edits and is used to evaluate the strength of an edit, whether unaffected neighboring relations remain unchanged, and the generation ability of the model post editing. The MultiCounterFact modification ensures there are no direct conflicts between edits in the dataset (i.e. no duplicate edits for a given subject and relation). We use a subset of 10,000 questions for our experiments.
@@ 4.2 Baselines
FT-W is a finetuning approach with weight decay to avoid altering other memories or changing unrelated model behavior (Zhu et al., 2020).
MEND uses a hypernetwork meta-trained to modify fine-tuning gradients so that they produce higher quality edits (Mitchell et al., 2021).
ROME uses counterfactual tracing to locate what layer of the model contains the information to be edited before updating the MLP component in that layer (Meng et al., 2022a).
MEMIT (presented above) uses a similar objective to ROME but modified to enable editing large sets of facts at once, which none of the other baselines are able to do well (Meng et al., 2022b). We consider this to be the main baseline as it is also the editing mechanism we use in our K-Edit implementation.
@@ 6 Related Works
These approaches include aforementioned works Zhu et al. (2020); Mitchell et al. (2021); Meng et al. (2022b, a). Other works use a hypernetwork to predict parameter updates for a model Cao et al. (2021)
@@
Outside of direct model editing approaches (Zhu et al., 2020; Meng et al., 2022a; Mitchell et al., 2021; Meng et al., 2022b; Ma et al., 2023), other approaches have sought to adapt to changing knowledge via storing external edit memories and employing various in-context learning and retrieval mechanisms. IKE and others add all edits in the LLM context (Zheng et al., 2023; Onoe et al., 2023); SERAC identifies if a related fact has been edited and retrieves it (Mitchell et al., 2022); MeLLo and DeepEdit iteratively retrieve edited memories during the generation and decoding process (Zhong et al., 2023; Wang et al., 2024).
We note that these retrieval-based editing approaches are fundamentally different than direct model editing approaches. Direct model editing leaves the model architecture and code entirely the same with the only change being to parameter weights. This means any system developed to use the original model can use the updated model with no other changes. In contrast, retrieval-based editing systems fundamentally change how the model is used and functions. For example, MeLLo requires setting up a retrieval database, iteratively prompting the LLM, making multiple calls to the database, and prompting the LLM to create further sub-questions Zhong et al. (2023). This fundamentally changes the characteristics of the model, increases latency, and alters model behavior for non-QA tasks. An alternate line of work aims to supplement the parametric knowledge of LLMs by augmenting the model prompt with information retrieved from an indexed corpus Lewis et al. (2020) or Knowledge Graphs Markowitz et al. (2024).
@@
Other papers have looked into other aspects of contextual knowledge for edits such as reversibility (Ma et al., 2023), multi-linguality (Wang et al., 2023c; Si et al., 2024), or other implied logical conditions given an edit (Wang et al., 2023a; Cohen et al., 2023; Li et al., 2023).
@@ Appendix C MEMIT
MEMIT (Meng et al., 2022b) works as follows. MEMIT determines a small change to the $L$-th layer activation of the last token of $s$ (e.g. [Kingdom] in the example above) such that $t_{r}(s)$ produces the desired output $o^{*}$. It then modifies a subset of the model weights to produce the desired activation for a whole batch of such edits.